SURFACE-ENHANCED PHOTOCHEMISTRY OF P-NITROBENZOIC ACID ON AG COATED AU COL-LOIDAL METAL FILMS AT 633 NM.

\underline{B. D. GILBERT}, Linfield College, Department of Chemistry, 900 SE. Baker St., McMinnville, Oregon 97128; J. MIRKOVIC, George Harrison Spectroscopy Laboratory, Massachusetts Institute of Technology, Cambridge, Massachusetts, 02139-4307; D. K. LAKER, 3614 Mandavilla Way, Apex, North Carolina 27502.

Surface enhanced Raman scattering (SERS) was used to probe the 633 nm photoreduction of p-nitrobenzoic acid (PNBA) to azodibenzoate on Ag and Ag coated Au colloidal metal films. The variation of the photolysis rate was determined by monitoring the growth of the 1470cm^{−1} N=N stretch of the photoproduct (azodibenzoate) compared with the 1600cm^{−1} aromatic quadrant stretch of the PNBA. The rate was determined from the ratio of the 1470cm^{−1} band to the 1600cm^{−1} band as a function of time. The rate was found to be much larger on bare Ag colloidal metal films than on Ag coated Au colloidal metal films, and to decrease as a function of the Ag coverage. These changes in the observed photoreduction rate are consistent with an electromagnetic enhancement mechanism.